# CORDIAL: Can Multimodal Large Language Models Effectively Understand Coherence Relationships?

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) are renowned for their superior instruction-following and reasoning capabilities across diverse problem domains. However, existing benchmarks primarily focus on assessing factual and logical correctness in downstream tasks, with limited emphasis on evaluating MLLMs’ ability to interpret pragmatic cues and intermodal relationships. To address this gap, we assess the competency of MLLMs in performing Multimodal Discourse Analysis (MDA) using Coherence Relations. Our benchmark, CORDIAL, encompasses a broad spectrum of Coherence Relations across 3 different discourse domains at varying levels of granularity. Through our experiments on 10+ MLLMs employing different prompting strategies, we show that even top models like Gemini 1.5 Pro and GPT-4o fail to match the performance of simple classifier-based baselines. This study emphasizes the need to move beyond similarity-based metrics and adopt a discourse-driven framework for evaluating MLLMs, providing a more nuanced assessment of their capabilities. The benchmark and code are available at: https://github.com/aashish2000/CORDIAL.
@@ 3.4 Baseline Classifier
Dataset CR Claude Gemini GPT4o
DisREL Similar 70.4% 57.2% 14.8%
Complementary 91.2% 10.8% 96.8%
Overall 80.8% 34.0% 55.8%
Tweet Subtitles Insertion 20.59% 0.0% 11.76%
Concretization 74.1% 57.35% 37.61%
Projection 81.82% 0.0% 15.91%
Restatement 65.73% 64.34% 21.68%
Extension 66.29% 0.0% 38.29%
Overall 70.44% 47.69% 34.56%
CLUE SL Visible 83.37% 90.21% 75.4%
Subjective 58.0% 20.0% 52.0%
Action 72.73% 9.09% 54.55%
Story 29.12% 3.85% 35.71%
Meta 9.98% 0.0% 0.8%
Overall 42.77% 35.0% 36.52%
CLUE ML Overall 48.82% 32.71% 44.21%
@@ 4.3 Prompting Strategies and Fine-tuning
In addition to zero-shot evaluation, we also investigate the contribution of few-shot and Chain-of-Thought (CoT) prompting strategies in enabling MLLMs to learn coherence relations better. For few-shot, we include one example per coherence relation in each prompt as examples in the 3 single-label classification settings. For multi-label classification on CLUE ML, we include 6 different examples covering different combinations of relations in our prompt. To perform CoT, we include a reasoning step in our prompt that asks the model to generate a rationale before predicting the coherence relation. More details about the prompt templates used for each of the tasks are present in Sections C.1 and D.1 of our appendix. We fine-tune the Llama 3.2 11B Instruct model on our benchmark to measure the impact of task-specifc fine-tuning in open-source MLLMs with hyperparameter selection described in Appendix Section E.
@@ 4.4 Main Results
From our results in Tables 2, 4, 5, 6 we observe that no MLLM shows improvements over our baseline classifier on Macro F1 scores across all settings. When strictly looking at zero-shot prompts, Claude 3.5 Sonnet v2 performs the best on Tweet Subtitles, CLUE ML, and CLUE SL while Gemini 1.5 Flash performs the best on DisRel. However, the CLIP Classifier can outperform these MLLMs by 2.4% on DisRel, 64.1% on Tweet Subtitles, 38.6% on CLUE SL, and 5.6% on CLUE ML in terms of Macro F1 score. This shows that although these datasets have clearly discernible visual and text features that help in predicting coherence relations, MLLMs aren’t able to comprehend them effectively. The trend extends to both proprietary and open-source MLLMs regardless of their size. Our results reiterate the need for benchmarks such as CORDIAL to evaluate the intermodal reasoning capabilities of MLLMs.
@@ 5 Conclusions
While our proposed benchmark provides a comprehensive assessment of intermodal reasoning in current MLLMs, several limitations must be acknowledged. Firstly, the benchmark is currently limited to analyzing coherence relations in single-turn discourses. This is due to a lack of publically available datasets that provide multi-turn image-text pairs with annotated coherence relations. We plan to extend our benchmark to include multi-turn discourse relations as future work. Secondly, although we analyze different discourse domains in our benchmark, we lack a unified set of coherence relations that can be applied across all domains. The difficulty in defining a universal set of coherence relations is due to the varying nature of discourse in different domains. This limits our ability to analyze the inter-domain performance of MLLMs on the same set of relations. Finally, our benchmark is currently limited to the English language and must be extended to multi-lingual discourses as well.
@@ C.1 Prompt Templates
  • User Message: This message is used to reiterate the task again, along with the required output format. The image and text that need to be evaluated, is also added here.
@@ Appendix E Fine-tuning Details
We fine-tune LLaMA 3.2 Vision 11B Instruct (unsloth/Llama-3.2-11B-Vision-Instruct in Huggingface) using the Unsloth [^3] https://unsloth.ai/blog/vision framework. We opted for this framework due to its memory efficiency and rapid fine-tuning. We perform Parameter Efficient Fine-Tuning (PEFT) of all layers (Vision & Language) and modules (Attention & MLP) present. We use the hyperparameters mentioned in Section E.1 on each dataset for fine-tuning. Other parameters have been initialized to their default values.
@@ Appendix G Computational Resources
Zero/Few Shot Prompt for DisREL, Tweet Subtitles and CLUE Single-Label System <insert-system-message> User Based on provided information, predict the most applicable Coherence Relation for the next image-text pair. Output only one relation (<insert-coherence-relations) and do not include any other information in your response. Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to finetuned LLaMA 3.2 Vision’s prompt in CLUE Single-Label, to enhance output format adherence.) <add-few-shot-examples> <insert-image-text-pair> Assistant Coherence Relation:
@@
Zero/Few Shot Prompt for CLUE Multi-Label System <insert-system-message> User Based on provided information, predict the correct Coherence Relations for the next image-text pair. Output them as a JSON value to the key labels" and do not include any other information in your response. (Default output format for all models) Give your predicted labels as comma separated values. Do not include any other information in your response. (Alternate output format for LLaMA 3.2, Phi 3.5, Qwen2-VL and LLaVA-OneVision) Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to LLaVA 1.6 13B prompt to enhance output format adherence.) <add-few-shot-examples> <insert-image-text-pair> Assistant Coherence Relations:
CoT Prompt for CLUE Multi-Label System <insert-system-message> User Before assigning a coherence relation, let’s think step by step and analyze the image-text pair in depth. <insert-image-text-pair> Assistant Analysis: <add-analysis-from-model> User Now, using your analysis, predict the correct Coherence Relations for the image-text pair. Output them as a JSON value to the key labels" and do not include any other information in your response. (Default output format for all models) Give your predicted labels as comma separated values. Do not include any other information in your response. (Alternate output format for LLaMA 3.2, Phi 3.5, Qwen2-VL and LLaVA OneVision) Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to LLaVA 1.6 13B prompt to enhance output format adherence.) Assistant Coherence Relations:

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) are renowned for their superior instruction-following and reasoning capabilities across diverse problem domains. However, existing benchmarks primarily focus on assessing factual and logical correctness in downstream tasks, with limited emphasis on evaluating MLLMs’ ability to interpret pragmatic cues and intermodal relationships. To address this gap, we assess the competency of MLLMs in performing Multimodal Discourse Analysis (MDA) using Coherence Relations. Our benchmark, CORDIAL, encompasses a broad spectrum of Coherence Relations across 3 different discourse domains at varying levels of granularity. Through our experiments on 10+ MLLMs employing different prompting strategies, we show that even top models like Gemini 1.5 Pro and GPT-4o fail to match the performance of simple classifier-based baselines. This study emphasizes the need to move beyond similarity-based metrics and adopt a discourse-driven framework for evaluating MLLMs, providing a more nuanced assessment of their capabilities. The benchmark and code are available at: https://aashish2000.github.io/CORDIAL/.
@@ 4 Experiments
Dataset CR Claude Gemini GPT4o
DisREL Similar 70.4% 57.2% 14.8%
Complementary 91.2% 10.8% 96.8%
Overall 80.8% 34.0% 55.8%
Tweet Subtitles Insertion 20.59% 0.0% 11.76%
Concretization 74.1% 57.35% 37.61%
Projection 81.82% 0.0% 15.91%
Restatement 65.73% 64.34% 21.68%
Extension 66.29% 0.0% 38.29%
Overall 70.44% 47.69% 34.56%
CLUE SL Visible 83.37% 90.21% 75.4%
Subjective 58.0% 20.0% 52.0%
Action 72.73% 9.09% 54.55%
Story 29.12% 3.85% 35.71%
Meta 9.98% 0.0% 0.8%
Overall 42.77% 35.0% 36.52%
CLUE ML Overall 48.82% 32.71% 44.21%
@@ 4.3 Prompting Strategies and Fine-tuning
In addition to zero-shot evaluation, we also investigate the contribution of few-shot and Chain-of-Thought (CoT) prompting strategies in enabling MLLMs to learn coherence relations better. For few-shot, we include one example per coherence relation in each prompt as examples in the 3 single-label classification settings. For multi-label classification on CLUE ML, we include 6 different examples covering different combinations of relations in our prompt. To perform CoT, we include a reasoning step in our prompt that asks the model to generate a rationale before predicting the coherence relation. More details about the prompt templates used for each of the tasks are present in Sections C.1 and D.1 of our appendix. We fine-tune the Llama 3.2 11B Instruct model on our benchmark to measure the impact of task-specific fine-tuning in open-source MLLMs with hyperparameter selection described in Appendix Section E.
5 Findings and Implications
@@ 5.1 Main Results
From our results in Tables 2, 3, 4, 6 we observe that no MLLM shows improvements over our baseline classifier on Macro F1 scores across all settings. When strictly looking at zero-shot prompts, Claude 3.5 Sonnet v2 performs the best on Tweet Subtitles, CLUE ML, and CLUE SL while Gemini 1.5 Flash performs the best on DisRel. However, the CLIP Classifier can outperform these MLLMs by 2.4% on DisRel, 64.1% on Tweet Subtitles, 38.6% on CLUE SL, and 5.6% on CLUE ML in terms of Macro F1 score. This shows that although these datasets have clearly discernible visual and text features that help in predicting coherence relations, MLLMs aren’t able to comprehend them effectively. The trend extends to both proprietary and open-source MLLMs regardless of their size. Our results reiterate the need for benchmarks such as CORDIAL to evaluate the intermodal reasoning capabilities of MLLMs.
@@ 5.2 Discussion
Cross-Discourse Generalization of CR Taxonomies
With our evaluation of MLLMs performing MDA, we show that they perform much worse compared to baseline classifiers within each discourse type. Since CR taxonomies are designed specifically for the discourses analyzed, it is difficult to directly infer the cross-domain translation of model performances from one discourse to another. Previous studies on text-only discourses report poor cross-domain adaptation of traditional classifiers across discourse types Bourgonje and Demberg (2024). As CORDIAL extends evaluation to multiple discourses, we are able to provide a better assessment of MLLM capabilities in Discourse Analysis. A natural extension of our benchmark would be designing a unified set of CRs that can be applied across complementary discourses. This setting would be especially challenging for our classifier baselines due to the varying distribution of images and text from different sources, compared to MLLMs which are more robust to these changes. The grouping of complementary discourse domains and definition of new unified CRs is a challenging task which we aim to investigate as a part of our future work.
@@ 6 Conclusions
While our proposed benchmark provides a comprehensive assessment of intermodal reasoning in current MLLMs, several limitations must be acknowledged. The benchmark is currently limited to analyzing coherence relations in single-turn discourses. This is due to a lack of publicly available datasets that provide multi-turn image-text pairs with annotated coherence relations. We plan to extend our benchmark to include multi-turn discourse relations as future work. Our benchmark is currently limited to the English language and must be extended to multi-lingual discourses as well.
@@ C.1 Prompt Templates
  • User Message: This message is used to reiterate the task again, along with the required output format. The image and text that needs to be evaluated, is also added here.
@@ Appendix E Fine-tuning Details
We fine-tune LLaMA 3.2 Vision 11B Instruct (unsloth/Llama-3.2-11B-Vision-Instruct in Huggingface) using the Unsloth [^3] https://unsloth.ai/blog/vision framework. We opted for this framework due to its memory efficiency and rapid fine-tuning capabilities. We perform Parameter Efficient Fine-Tuning (PEFT) of all layers (Vision & Language) and modules (Attention & MLP) present. We use the hyperparameters mentioned in Section E.1 on each dataset for fine-tuning. Other parameters have been initialized to their default values.
@@ Appendix G Computational Resources
Zero/Few Shot Prompt for DisREL, Tweet Subtitles and CLUE Single-Label System <insert-system-message> User Based on provided information, predict the most applicable Coherence Relation for the next image-text pair. Output only one relation (<insert-coherence-relations>) and do not include any other information in your response. Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to finetuned LLaMA 3.2 Vision’s prompt in CLUE Single-Label, to enhance output format adherence.) <add-few-shot-examples> <insert-image-text-pair> Assistant Coherence Relation:
@@
Zero/Few Shot Prompt for CLUE Multi-Label System <insert-system-message> User Based on provided information, predict the correct Coherence Relations for the next image-text pair. Output them as a JSON value to the key "labels" and do not include any other information in your response. (Default output format for all models) Give your predicted labels as comma separated values. Do not include any other information in your response. (Alternate output format for LLaMA 3.2, Phi 3.5, Qwen2-VL and LLaVA-OneVision) Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to LLaVA 1.6 13B prompt to enhance output format adherence.) <add-few-shot-examples> <insert-image-text-pair> Assistant Coherence Relations:
CoT Prompt for CLUE Multi-Label System <insert-system-message> User Before assigning a coherence relation, let’s think step by step and analyze the image-text pair in depth. <insert-image-text-pair> Assistant Analysis: <add-analysis-from-model> User Now, using your analysis, predict the correct Coherence Relations for the image-text pair. Output them as a JSON value to the key "labels" and do not include any other information in your response. (Default output format for all models) Give your predicted labels as comma separated values. Do not include any other information in your response. (Alternate output format for LLaMA 3.2, Phi 3.5, Qwen2-VL and LLaVA OneVision) Use the format "Coherence Relation: <insert-coherence-relation>" for your response. (Added to LLaVA 1.6 13B prompt to enhance output format adherence.) Assistant Coherence Relations: